# Does differentially private synthetic data lead to synthetic discoveries?

### v1

@@ 6 Discussion
This study investigated to what extent the validity and power of the MW U test are preserved in DP-synthetic data. The experimental results on both Gaussian and real-world data demonstrate, that DP-synthetic data, especially with strong privacy guarantees ($\epsilon\leq 1$) can lead to false discoveries, specifically in the case of the MW U test. We empirically show that many state-of-the-art DP methods for generating synthetic data have highly inflated Type I error when the privacy level is high. These results indicate, that false discoveries or inferences are likely to be drawn from the DP-synthetic data produced by these DP methods. Our findings are in line with other studies that have presented or stated that DP-synthetic data can be invalid for statistical inference and indicated the need for methods that are noise-aware in order to produce accurate statistical inferences [17, 54, 55, 56, 57].
@@
There are several limitations in our study that could be addressed in future research. One limitation is that we generated datasets with only two variables, which is a very simple case. However, even with this simple case, we empirically showed the risk of false inference when using DP-synthetic data. Increasing the dimensionality of the dataset will increase the complexity of the generation process, which may result in a higher risk of false inference. Another limitation is that marginals or histograms based DP methods require continuous variables to be discretized. This discretization must be performed in a private matter or based on literature to avoid leaking private information. Besides, it is well-known that the number of bins used to discretize the data has a significant impact on the quality of the resulting data [16, 58]. Therefore, choosing the number of bins is problem- and data-dependent and can affect the results. In our experiments with Gaussian data, the continuous values were discretized using 100 bins. This number of bins was selected to show a possible extreme case where having bins empty or with small counts deteriorates the quality of the generated DP-synthetic data. On the other hand, for our experiments with real-world data, the number of bins used was based on literature. Finally, testing different hyperparameter values for the DP method implementations could yield different results for the methods.
@@ 7 Conclusions
Our results suggest caution when releasing DP-synthetic data, as false discoveries or loss of information are likely to happen especially when a high level of privacy is enforced. To an extent, these issues may be mitigated by having large enough original datasets, selecting methods that are less prone to adding false signal to data, and by carefully comparing the quality of the DP-synthetic data to the original one based on various quality metrics (see e.g. [4]) before data release. Still, with current methods DP-synthetic data may be a poor substitute for real data when performing statistical hypothesis testing, as one cannot be sure if the results obtained are based on trends that hold true in the real data, or due to artefacts introduced when synthetizing the data.
@@ 8 Acknowledgements
This work has received funding from Business Finland (grant number 37428/31/2020) and European Union’s Horizon Europe research and innovation programme (grant number 101095384).The authors would like to express their gratitude to Peter J. Boström, Ivan Jambor, and collaborators for their support and contribution in providing the prostate cancer datasets used in the real-world data experiments.

### v2

@@ 6 Discussion
This study investigated to what extent the validity and power of independent sample tests are preserved in DP-synthetic data. Experimental results on Gaussian, real-world, and multivariate simulated data demonstrate that the generated DP-synthetic data, especially with strong privacy guarantees ($\epsilon\leq 1$), can lead to false discoveries. We empirically show that many state-of-the-art DP methods for generating synthetic data have highly inflated Type I error rates when the privacy level is high. These results indicate that false discoveries or inferences are likely to be drawn from the DP-synthetic data produced by these DP methods. Our findings are in line with other studies that have presented or stated that DP-synthetic data can be invalid for statistical inference and indicated the need for methods that are noise-aware in order to produce accurate statistical inferences [17, 66, 67, 68, 69].
@@
The main advantage of releasing DP-synthetic data, as opposed to releasing only analysis results from the original data, is that it can be ideally used to support a wide range of analyses by different users. Due to post-processing property of DP, any type or number of analyses done on the synthetic data are also guaranteed to be DP with no further privacy budget needed. However, if the only goal is to perform a limited number of pre-defined analyses, it makes more sense to do these on the original data with DP methods. This is illustrated in our experiments by the DP-MW U test baseline that always outperforms analyses done on DP-synthetic data. As a middle ground between these approaches, an active area of research is to develop such DP synthetization methods where the data is optimized to support certain types of analyses well, such as PrivPfC for classifier training [70] and various Bayesian noise-aware DP synthetic data generation methods [69].
There are limitations in our study that could be addressed in future research. One limitation is that the marginals or histograms based DP methods require continuous variables to be discretized. This discretization must be performed in a private matter or based on literature to avoid leaking private information. Besides, it is well-known that the number of bins used to discretize the data has a significant impact on the quality of the resulting data [16, 43]. Therefore, choosing the number of bins is problem- and data-dependent and can affect the results. In our experiments with Gaussian data, the continuous values were discretized using 100 bins. This number of bins was selected to show a possible extreme case where having bins empty or with small counts deteriorates the quality of the generated DP-synthetic data. On the other hand, for our experiments with real-world and multivariate simulated data, the number of bins used was determined based on domain knowledge and literature. Finally, testing different hyperparameter values for the DP method implementations could yield different results for the methods.
@@ 7 Conclusions
Our results suggest caution when releasing DP-synthetic data, as false discoveries or loss of information are likely to happen especially when a high level of privacy is enforced. To an extent, these issues may be mitigated by having large enough original datasets, selecting methods that are less prone to adding false signal to data, and by carefully comparing the quality of the DP-synthetic data to the original one based on various quality metrics (see e.g. [4]) before data release. Still, with current methods DP-synthetic data may be a poor substitute for real data when performing statistical hypothesis testing, as one cannot be sure if the results obtained are based on trends that hold true in the real data, or due to artefacts introduced when synthesizing the data.
@@ 8 Acknowledgements
This work has received funding from Business Finland (grant number 37428/31/2020) and European Union’s Horizon Europe research and innovation programme (grant number 101095384). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. The authors would like to express their gratitude to Peter J. Boström, Ivan Jambor, and collaborators for their support and contribution in providing the prostate cancer datasets used in the real-world data experiments. We also thank Katariina Perkonoja for her insightful feedback regarding the experimental setup and the statistical tests, as well as the anonymous reviewers for their valuable comments.
@@ A.1 Drawing data from additively smoothed histograms is differentially private
A.2 Failed DP-synthetic data sets
DP-synthetic data generation methods can sometimes produce data sets that fail to meet the requirements of a statistical test, even if the original sensitive data would fulfill these. This type of degenerate behavior was observed in the experiments especially when generating multivariate data for very small $\epsilon$ values and small dataset sizes. Typical failure mode observed was that only one class was represented in the synthetic data, in which case none of the tests apply. Further, in some cases all the values of continuous variable were observed to be the same in the synthetic data, in which case t-test is undefined. In the case of the chi-squared test applied to binary variables, highly imbalanced or sparse synthetic data resulted in low expected frequencies, meaning that one or more cells in the contingency table had an expected frequency of less than five, causing the test to fail. In the experiments, Type I and Type II was calculated based on only those generated data sets that met the requirements of the considered tests. Tables S1 and S2 catalogue the number of failed data sets for the experiments performed on the multivariate simulated prostate cancer data.
Table S1: Number DP-synthetic datasets that failed to meet the requirements of the statistical test out of 1000 generated for each corresponding DP-synthesizer, privacy budget, and dataset size on simulated multivariate non-signal data. There were no failures in experiments with sample sizes of 20 000, as well as in datasets with sizes of 50, 100, 500, and 100 and $\varepsilon\geq 5$.
Table S2: Number DP-synthetic datasets that failed to meet the requirements of the statistical test out of 1000 generated for each corresponding DP-synthesizer, privacy budget, and dataset size on simulated multivariate signal data. There were no failures in experiments with sample sizes of 20 000, as well as in datasets with sizes of 50, 100, 500, and 100 and $\varepsilon\geq 5$.